LUCIFER Training & Preparations for Partner Science:
UT 2010 May 04

Observer: DThompson (LBTO)
Telescope Operator: WWack (Tucson)
Instrument Support: TShih
Telescope Support: JLittle, MMidkiff
SW support: SHooper (on call)

Summary

  • Previous night (20100503) lost to bad weather, no log kept

  • Twilight flats in Br_gam, H2, FeII, and a dithered set in Ks on "blank" field.
  • Ks imaging on two B3.2 radio galaxies
  • Ks imaging on CSS100217 transient...it looks like a point source to me
  • OSU source: SDSS J0942+5207 J spectroscopy + calibrations
  • OSU source: UM425 J spectroscopy + calibrations
  • MOS error required Peter's intervention, 90m lost
  • INAF: 1 hour K spectroscopy on HDF
  • LBTO: Test spectrum on XRay Binary
  • LBTO: Ks imaging on HR10 to end of the night.

Notes for LUCIFER Team:
  • Formatting issue on guide star coords (source DEC reads +00 04 37.6, GS reads 0:7:53.71)
  • Possible detector focus stage issues? Poor imaging on LUCIFER but good on guider, corrected by initializing LUCIFER.

Details

00:10 Clear outside. Tout=4.6 Tmirror=-0.9 RH=29% Dewpt=-11.7 wind up to 10m/s. Sunset is at UT02:12. LUCIFER data: luci.20100504.NNNN.fits

01:50 Twilight flats (2s o2dcr):
  • 01:50 BrG Files: 0002-0011 11kDN
  • 01:54 H2 Files: 0012-0021 12kDN
  • 02:04 BrG Files: 0022-0031 6kDN
  • 02:06 H2 Files: 0032-0041 6kDN
  • 02:08 FeII Files: 0042-0051 8kDN
  • 02:10 FeII Files: 0052-0061 6kDN

02:12 Pointing correction, IE=-19.6, CA=4.8, IA=-393 (from last night)
  • WT10_305 0064 EL~84.6 IE now -34 CA now -12
  • ACT0084 0066 EL~21 IA now -396
  • WT10_305 0067 EL~8 IE now -36 CA now -9

02:27 Twilight flats in Ks, TRACK mode at 09:44:30 00:10:00, dithered
  • Files 0068 - 0069 only 700 DN in 2 seconds!
  • Files 0070 - 0080 ~3kDN in 10 seconds

02:33 Preset to BS9138 for collimation, on-axis. GCS Error: couldn't request a full CCD readout for GS acquisition. Error is -90. Repeated the same error 3 times. Calling Torsten (02:36) - needed to restart GCS because the AGw was restarted earlier.

02:37 Preset to BS9138 for collimation on-axis. Clear temp-induced offsets in WFS images.

LBTO: B3.2USS radio galaxies
02:42 Preset to B3.2 0902_0021. 02:44 execute Ks_600s script. Files 0081-0090. FWHM~0.66 arcsec in Ks.

03:00 Preset to B3.2 0947+0004. 03:03 execute Ks_600s script. Files 0091-0100. FWHM~0.9 arcsec in Ks.

LBTO CSS100217+404220 Transient 03:19 Preset to CSS100217+404220 by executing Ks_imaging script. Odd, the guide star was found but it jumped out of the guide image after guiding was started. Was this a jump? EL~80.

03:24 Preset to CSS100217 again. Files 0101-0106. Source clearly have odd distortion, doubled. Aborting script after 6 images, clearing all active optics etc. IE=-28, CA=-14.

03:46 going back to CSS100217, will collimate on guide star. WFS still having trouble converging. Files 0109- Turned OFF left M1 ventilation around file 0120. Tried moving to N1.8 and back and there was an immediate improvement. LUCIFER now <0.5 arcsec. Could the detector focus stage have been contributing to the IQ problems? At least some of the images should be usable. CSS looks like a point source.

OSU: SDSS J0942+5207
04:36 Preset to OSU source: SDSS J0942+5207 [04:04:38.41 Error while moving mask on logging window, but the mask move appears to have completed OK].
  • File 0139 through slit image
  • File 0140 on-sky image (offset = -0.81, 0.91
  • File 0141 source in through-slit image
    • DIQ not good...source is obviously elongated
    • FWHM=0.45 arcsec on guider, but also distorted

04:48 Started spectroscopic science script, Files 0142 - 0147. Looks good.
05:16 Started ff calibration script after moving in calib. unit.
05:20 Started wave calib script. Last image 0167

05:25 Correct pointing near Telluric...0168 IE=-14 CA=-26

05:28 Started script for HIP50303 telluric standard
  • 0169 through slit
  • 0170 img
  • 0171 through slit with Pa_beta + J to see slit
  • 0172 source in slit.

05:39 Started script for HIP50303 science, files 0173-0176

OSU: UM425
05:45 Started UM425 acq, files 177-179
05:56 Started UM425 sci, files 0180-0185

06:22 Started HIP54849J_acq, files 0186-0189
06:32 Started HIP54849J_sci, files 0190-0195

06:39 Started ff calibration script @ telluric position, files 0196-205
06:43 Started wave calib script, files 0206-

INAF: HDF Mannucci
06:49 Pointing correction near HDF, file 0214. IE now -27, CA now -34.

06:58:22 Error in sequence MOS - storage to turnout. WARNING holding current still active at [card number 1, motor number 5, type: "mask handling unit head translation". LUCIFER Team emailed, Walter called. Peter is in transit and will look into the problem when he gets into work at 8:00UT. Fixed at 8:31, thanks Peter!

06:49 Pointing correction near HDF, file 0215. IE now -26, CA now -33.

08:37 mannucci_acq_HDF script. Guide star is a close double star (should be OK).
  • 0216 sky image
  • 0217 source image
  • 0218 through-slit image (on sky)
  • 0219 source on slit image. All 3 sources that should be visible are...

09:01 Started K_3_PRIORITY1 script. Note LST is 16:31, EL=43 at start. Files 0220-231. Field is setting, but the seeing is good and the sky is clear. Sources on LUCIFER clearly elongated. Mirror now in good thermal equilibrium. By file #227 the guider images were obviously elongated, ~1 arcsec and up to ~3. We are facing close to into the 10m/s wind, and the guide star is occasionally round, so this is likely persistent vibrations.

10:01, just before the last spectrum on the HDF mask, LUCIFER started sending many flexure compensation errors. Telescope is at EL~36, and ROT~295.

10:08 Started flat script in K with the calibration unit in.
10:12 Started arc script in K. Files through 0241

10:20 Started telluric acquisition for HDF. Failed...no guide star. Correcting pointing...N3.75 camera out of position (switch to N1.8 and back manually, position is better now): IE now -45 CA now -42. Changed telluric acq to have mask_in_turnout first as that is where it will end up for the pointing correction.

10:27 Started telluric acquisition for HDF. File 0243 is star image, 0244 is through-slit image (star there), 0245 is after +0.36 arcsec offset in X.

10:33 Started telluric science script for K. Files 0246 - 0253 Elevation ~34 degrees. Looks out of focus.

LBTO: XRBinaries 10:36 Started Slit07_acq script.
  • 0254 Sky image
  • 0255 source image
  • 0256 slit image
  • 0257 source in slit image

10:53 Started Slit07_sci script. EL~27 degrees, guider showing 0.8 arcsec FWHM, reasonable collimation. Files 0258 - 0259 Also ran specflatK (files 0260-0263)and specarcK (files 0264-) scripts with calib unit. HIP88566 A0V V=8.32 mag for telluric. 270 image, 271 slit, 272 source+slit, 273 and 274 spectra.

LBTO: HR10
11:24 There is only ~35 minutes left, not enough time for a spectrum, check DIQ - not good, everything is elongated and the guider looks OK. Initialized LUCIFER and the images look good now. Guider seeing 0.45 arcsec FWHM.

11:55 Done...closing dome.

12:07 Started DoDarks.sh script.

-- DavidThompson - 04 May 2010
Topic revision: r8 - 05 May 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback